# Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models

### v1

@@ 3.3 Post-processing and Unified Task format
Table 2: General RAG citation quality and answer correctness results on ExpertQA.
4 Model Training
The general-purpose LLM autoevaluator that we trained is available in two sizes, and they are denoted as REC-12B and REC-70B separately. REC-12B is instruction fine-tuned from Mistral-Nemo [^3] https://mistral.ai/news/mistral-nemo, and REC-70B is instruction fine-tuned from Llama-3.1-70B Team (2024). We adopted supervised fine-tuning (SFT) to optimize the models. To accommodate GPU memory constraints, we made the following design choices: we filtered out examples where the combined length of the prompt and the response exceeded 6,144 tokens. We used Low-Rank Adaptation (LoRA) during fine-tuning Hu et al. (2021), with a rank of $r=256$ for REC-12B and $r=64$ for REC-70B. Additionally, We applied 4-bit quantization for REC-70B during initialization. Both models were trained using a learning rate of $1\times 10^{-4}$, with data shuffled and for a single epoch. REC-12B was trained with a batch size of 2 and a gradient accumulation factor of 8, whereas for REC-70B we used a batch size of 1 with the same gradient accumulation factor. Both models were trained on eight NVIDIA H100 GPUs, each with 80GB of memory. The total training time was approximately 4.5 hours for REC-12B and 5.5 hours for REC-70B.
5 Evaluation
In this section, we evaluate our REC models on diverse benchmarks assessing LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), each of which is an important measure to a general-purpose LLM autoevaluator. We compare our results against 7 SOTA LLMs, including Misral-7B (Mistral-7B-Instruct-v0.2), Mistral-Nemo (Mistral-Nemo-Instruct-2407), Llama-3.1-70B (Llama-3.1-70B), Claude-3-Opus (claude-3-opus-20240229), GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-turbo), GPT-4o (gpt-4o).
@@ 5.1 RAG citation & Correctness
The ALCE (Automatic LLMs’ Citation Evaluation) benchmark is designed to assess the ability of LLMs to generate text with accurate and relevant citations. ALCE addresses this by providing a framework that evaluates the quality of citations in text generated by LLMs, focusing on three key dimensions: fluency, correctness, and citation quality. The ALCE benchmark is built on three datasets: ASQA (a short-answer question dataset), QAMPARI (which provides lists of correct answers), and ELI5 (a long-form question-answer dataset). For each dataset, ALCE evaluates how well the model generates text that is not only fluent and accurate but also properly supported by citations. The evaluation includes metrics like citation precision (ensuring all cited sources are relevant) and citation recall (ensuring all necessary sources are cited). ALCE does not provide training data but instead measures citation performance through retrieval-based systems that simulate real-world information-seeking tasks, where the model retrieves relevant passages and uses them to support its answers.
As shown in Table 1, our REC models rank highly in terms of overall average performance, with REC-70B achieving an average score of 41.62, the highest among compared models, and REC-12B reaching 39.95. This highlights the robustness and balance of these models across fluency, correctness, and citation quality. REC offers a well-rounded solution across different types of questions, making it highly reliable for real-world applications where both citation quality and answer correctness are essential. Specifically, while GPT-4 excels in citation quality, the much smaller REC models outperform all the other competitors.
@@
Rating accuracy Explanation accuracy Citation (F1) Rating accuracy Explanation accuracy Citation (F1) Rating accuracy Explanation accuracy Citation (F1)
@@
Table 3: Human evaluation results on ABCD summarization content quality evaluation task. For the correctness of rating and explanation, we report the average results of the two labelers. For citations, we evaluate the intersection of the results from both labelers against the machine-generated citations to compute F1.
ExpertQA
ExpertQA Malaviya et al. (2024) is designed to evaluate LLMs’ ability to generate accurate and well-attributed responses in technical and high-stakes fields, such as medicine and law. It was developed by involving experts from 32 different fields, who contributed 2,177 domain-specific questions based on their knowledge. In total, 484 participants helped curate these questions. LLMs were then used to generate responses to the questions, followed by human experts evaluating the quality of these responses on several criteria, including factual correctness, completeness of attribution, source reliability, and informativeness. We adopt the metrics used in its paper: AutoAIS Gao et al. (2022), which is similar to citation recall, and FActscore Min et al. (2023) which measures the percentage of generated claims that are factual. We generate claims from gold long-form answers with every compared model, and use an NLI model Honovich et al. (2022) to determine the attribution of claim-evidence pairs in the ExpertQA dataset.
As shown by the zero-shot results in Table 2, our REC-12B and REC-70B achieve the highest AutoAIS and FActscore scores, which not only indicates their best citation quality but also the generalizability of their attribution capability across domains.
@@ 5.2 Rating & Explanation & Citation
The ABCD dataset Chen et al. (2021) contains customer support conversation transcripts. We generate summaries by prompting Mistral-7B (Jiang et al., 2023) with: “Summarize knowledge from transcripts after they’ve ended, including the customer issue and resolution." to form (dialogue, summary) pairs for evaluation. LLM autoevaluators then take as input the summarization task prompt hydrated with the dialogue, and the summary, to generate rating, explanation, and citations according to metrics defined in Section 1. Note that we use a held-out set of 50 examples in this experiment to make sure there exists no data overlap with the ABCD subsets used during training. However, unlike all the other evaluations included in this paper, this evaluation is in-distribution.
Different from the general RAG citation benchmarks evaluated in Section 5.1, which contain human annotated citation groundtruth, there exists no well-established benchmarks to evaluate content quality citations. Hence, we seek human annotations to serve as a reference standard in this task. The final citation outputs were labeled by 12 machine learning experts where each has a graduate degree in computer science or related fields. They were asked to complete three tasks: rating correctness evaluation, explanation correctness evaluation, and to provide manually-written citations for the claims in explanation (See Appendix D for human labeling guideline).
We assign two labelers for each model and find the average inter-rater agreement is around 95%. For the correctness of rating and explanation, we report average results of the two labels, while for citations, we evaluate the intersection of the results of the two labelers against the machine-generated citations to compute F1 score as shown in Table 3. Due to the high quality summary generated by the summarizer LLM, most LLM autoevaluators obtain high rating and explanation accuracy. As for the citation task, both REC-12B and REC-70B models consistently outperform other models in F1 score.
@@
LLM-FactVerify Wiki-FactVerify Summarization Long-form QA AVERAGE
Mistral-7B 68.04 56.82 55.80 62.80 60.86
Mistral-Nemo 70.18 65.27 62.14 60.88 64.62
Llama-3.1-70B 75.25 68.84 69.56 76.98 72.66
Claude-3-Opus 76.82 76.52 70.65 76.82 75.20
GPT-3.5 72.81 69.53 68.60 70.40 70.33
GPT-4 77.44 77.22 72.45 76.92 76.01
GPT-4o 76.31 79.48 72.18 75.22 75.80
REC-12B 78.78 85.34 72.51 77.23 78.46
REC-70B 79.67 87.48 74.04 79.12 80.07
Table 5: LLM-AggreFact performance across four common use-cases: LLM-FactVerify (ClaimVerify + FactCheck + Reveal), Wiki-FactVerify (WiCE), Summarization (AggreFact + TofuEval), and Long-form QA (ExpertQA + LFQA). REC-70B outperforms all tested LLM-as-a-Judge models in all four categories.
@@ 5.3 General Capabilities
The RewardBench dataset assesses LLMs’ performance across several critical abilities using curated chosen-rejected response pairs Lambert et al. (2024). The evaluation process involves determining the model’s win percentage based on its ability to score the “chosen" response higher than the “rejected" one. Specifically, it’s focused on testing chat performance, safety measures, and reasoning (both in terms of code and math skills), while controlling for potential biases such as overfitting to prior datasets.
First we create prompts following the RewardBench evaluation format, where each prompt consisting of a query and two potential responses. The task for LLM is to evaluate these responses and determine which one is better. The performance of the reward models is quantified by calculating the win percentage for chosen-rejected pairs associated with each prompt. A “win" is defined by a scenario where the LLM presents a preference for the selected answer. This quantitative measure provides a clear benchmark for evaluating model performance across tasks. By averaging the results across the 4 core sections: Chat, Chat Hard, Safety, and Reasoning, a comprehensive measure of each model’s win percentage is derived. Our REC-12B achieves SOTA performance among all generative LLM with fewer than 20 billion parameters on the RewardBench leaderboard by October 15th, 2024, while our REC-70B achieves SOTA among all generative LLMs within the RewardBench Leaderboard by October 15th, 2024 as shown in Table 4.
@@
This is a benchmark for measuring the grounding capabilities of autoevaluators. Given a reference document and a claim, the autoevaluator determines if the claim is fully supported by the document. This holistic benchmark combines 10 attribution datasets used in recent studies on LLM factuality. Table 5 presents our attribution results on LLM-AggreFact Tang et al. (2024b), categorized into four common use-cases: (1) LLM-FactVerify: fact verification of LLM-generated responses, (2) Wiki-FactVerify: evaluating correctness of Wikipedia claims, (3) Summarization: assessing faithfulness of summaries, and (4) Long-form QA: evaluating long-form answers to questions. Our proposed model REC-70B outperforms all other models in all four categories. REC-70B achieves the highest overall average performance of 80.07, while the next-best baseline model REC-12B obtains a score of 78.46. In long-form QA attribution evaluation, our REC-70B model outperforms GPT-4 (80.07 vs 76.01), demonstrating its strong performance across all categories.
Given some of the models like GPT-3.5 have lower context windows, we removed examples longer than 16K tiktoken tokens from our evaluation. we use the same prompt instructions across categories, as shown in Appendix E. To reduce model API costs, we randomly sample 256 examples per evaluation task similar to the approach used in the FLAMe Vu et al. (2024b).
@@ 5.4 Bias Testing
Table 6 presents the performance of the compared models on the CoBBLEr benchmark with metrics of order bias, bandwagon effect, compassion, selective bias, salience, distraction, and frequency. Note that lower scores indicates better performance  (i.e., fewer biases) on CoBBLEr. Overall, our REC-70B model has the best average performance with a score of 0.2141, followed closely by GPT-4 (0.2279) and GPT-4o (0.2349). REC-12B, on the other hand, outperforms all the other same-sized and smaller models (Mistral-Nemo and Mistral-7B), and even some larger models (Llama-3-70B and GPT-3.5). The results suggests that training with REC leads to fewer cognitive biases and more consistent evaluation capabilities. In contrast, off-the-shelf LLMs can be more influenced by factors such as the order of responses or the length of the text.
@@ 6 Conclusions
LLMs excel in generating coherent and high-quality text for various tasks, but evaluating content quality is challenging due to issues like factual inaccuracies and hallucinations. This paper introduces two fine-tuned models—REC-12B and REC-70B—designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, completeness. These models not only provide ratings for these metrics but also offer detailed explanations and verifiable citations, i.e., Ratings, Explanations, and Citations (REC), thereby enhancing trust in the content. To achieve this goal, a curated dataset for explanations and citations fine-tuning is proposed to facilitate future research, which is the first public dataset containing both content quality citations and RAG citations. The models support various citation modes, including Post-fix citation, Post-fix citation mode with snippet, Inline citation, and Inline citation mode with snippet to balance the trade-off between latency and granularity of citation. Extensive evaluations on diverse benchmarks—ALCE, ExpertQA, ABCD summarization, RewardBench, LLM-AggreFact, and CoBBLEr—demonstrate that the fine-tuned LLM autoevaluator, REC-70B, outperforms state-of-the-art models. It excels in content evaluation, delivering clear explanations and ensuring citation accuracy.
@@
Our work introduces the innovative Rate, Explain, and Cite (REC) model for trusted automated AI evaluations, offering significant benefits but with a few limitations that future research will address. First, our work is limited to the baselines we compare, as non-LLM citation methods, such as semantic similarity, were not included. Including such techniques could serve as valuable baselines to further validate our model’s generalization and effectiveness. Second, the need for post-generation explanations and citations may introduce latency. Future work will compare latencies across models and explore optimizations to improve speed. We are also considering fine-tuning the snippet output using starting and ending character numbers to optimize for latency. Finally, even though the 12B base model is capable of generating multi-lingual output, our current evaluations are limited to English datasets. Expanding to multilingual evaluations is a key goal for future research.
@@
Our contributions present a novel general-purpose autoevaluator that supports various modes of citations, as well as a dataset with citations and explanations to support further research. There are a few ethical considerations for users of such technology. As with the use of fine-tuning and RAG, it is important to be aware of potential propagating biases and feedback loops. Biased data used for training and/or retrieval could lead to model outputs that reinforces such biases, impacting model quality and fairness. Although we extensively evaluated our LLM autoevaluator and found it to be safe, unbiased, and factual, users should consider whether such considerations hold true for the input source itself. Of note, even established sources such as knowledge banks may contain inaccurate or inconsistent information. Finally, though we aimed to be transparent and consistent for our data labeling tasks, using clear instructions and validations, note that data labeling tasks are inherently subject to the bias of labeler background and experiences.

### v2

@@ 6 Evaluation
Table 1: General RAG citation quality and answer correctness results on ExpertQA. The best and the second best performance for each metric are in bold and underlined separately.
Fluency Correct Citation F1
Mistral-7B 64.49 15.12 42.94
Mistral-Nemo 61.35 18.03 41.32
Llama-3.1-70B 72.01 19.09 40.18
Claude-3-Opus 67.53 12.11 35.30
GPT-3.5 51.20 19.02 50.87
GPT-4 57.98 20.08 47.09
GPT-4o 56.29 22.47 40.50
REC-8B 68.20 18.01 38.77
REC-12B 64.94 19.60 43.63
REC-70B 73.08 20.52 41.73
Table 2: Average fluency, answer correctness and citation quality across datasets on ALCE. The best and the second best performance for each metric are in bold and underlined separately.
@@ 6.1 RAG citation & Correctness
The two benchmarks discussed in this section evaluate LLMs’ ability to generate RAG citation in the inline mode.
ExpertQA
ExpertQA Malaviya et al. (2024) is designed to evaluate LLMs’ ability to generate accurate and well-attributed responses in technical and high-stakes fields, such as medicine and law. It was developed by involving experts from 32 different fields, who contributed 2,177 domain-specific questions based on their knowledge. In total, 484 participants helped curate these questions. LLMs were then used to generate responses to the questions, followed by human experts evaluating the quality of these responses on several criteria, including factual correctness, completeness of attribution, source reliability, and informativeness. We adopt the metrics used in its paper: AutoAIS Gao et al. (2022), which is similar to citation recall, and FActscore Min et al. (2023) which measures the percentage of generated claims that are factual.
As shown by the zero-shot results in Table 1, our REC-12B and REC-70B achieve the highest AutoAIS and FActscore scores, and our REC-8B is also able to outperform other similar-sized models (Mistral-7B and Mistral-Nemo), which not only indicates their better citation quality but also the generalizability of their attribution capability across domains.
@@
The ALCE (Automatic LLMs’ Citation Evaluation) benchmark (Gao et al., 2023) is designed to assess the ability of LLMs to generate text with accurate and relevant citation, focusing on three key dimensions: fluency, answer correctness, and citation quality. The ALCE benchmark is built on three datasets: ASQA (a short-answer question dataset), QAMPARI (which provides lists of correct answers), and ELI5 (a long-form question-answer dataset). Note that fluency is not measured in QAMPARI because of the nature of its answers being in lists. For each dataset, ALCE evaluates how well a model generates text that is not only fluent and accurate but also properly supported by citation. The citation quality evaluation includes citation precision (ensuring all cited sources are relevant) and citation recall (ensuring all necessary sources are cited).  [^8] For a detailed definition of the fluency and answer correctness metrics, please refer to Gao et al. (2023).
From Table 2, we see that our REC-70B model ranks highly in fluency and answer correctness. Although the GPT models excel in citation quality in these datasets, the much smaller REC-12B and REC-70B models outperform most of the competitors. When taking the three metrics together and considering the overall average performance, REC-70B achieves an average score of 45.11, the highest among the compared models. Since all criteria are important in generating a good response with accurate citations  [^9] For why the three metrics are all important to a verifiable LLM generation, see Appendix E., this highlights the robustness and balance of REC-70B across fluency, correctness, and citation quality. REC offers a well-rounded solution for different types of questions, making it highly reliable for real-world applications where both citation quality and answer correctness are essential.
@@
Rate Acc Explain Acc Citation (F1) Rate Acc Explain Acc Citation (F1) Rate Acc Explain Acc Citation (F1)
@@
Table 3: Human evaluation results on ABCD summarization content quality evaluation task.
@@ 6.2 Rating & Explanation & Citation
The benchmark discussed in this section evaluates LLMs’ ability to generate content quality citation in the inline with context snippet mode.
@@
The ABCD dataset Chen et al. (2021) contains customer support conversation transcripts. We generate summaries by prompting an open-source LLM, Mistral-7B (Jiang et al., 2023b), with: “Summarize knowledge from transcripts after they’ve ended, including the customer issue and resolution.” to form (dialogue, summary) pairs for evaluation. LLM autoevaluators then take as input the summarization task prompt, and a (dialogue, summary) pair, to generate rating, explanation, and citation according to the metrics defined in Section 1. Note that we use a held-out set of 50 examples in this experiment so there exists no data overlap with the ABCD subsets used during training.
Unlike the general RAG citation benchmarks evaluated in Section 6.1, which contain human annotated citation groundtruth, there exists no well-established benchmarks to evaluate content quality citation. Hence, we seek human annotations to serve as a reference standard in this task. The machine outputs were labeled by 12 machine learning experts, each of whom has a graduate degree in computer science or related fields. They were asked to complete three tasks: rating correctness evaluation, explanation correctness evaluation, and citation. For rating and explanation evaluation, the annotators were asked to provide “Yes/No” labels to determine the correctness of the responses generated by the models. For citation evaluation, they were asked to to provide manually-written citation for the claims in the generated explanation to serve as the groudtruth citation. (Details of the human labeling guideline can be found in Appendix F). Human evaluation on the Coherence metric was not performed because it is conceptually challenging to define what should be cited in a coherent response (e.g., citing an entire paragraph adds little interpretive value). Citation for coherence are more meaningful in negative cases where parts of the response are incoherent. To ensure a balanced evaluation with both positive and negative cases across metrics, we decided to exclude Coherence from the ABCD summarization content quality task.
We assign two labelers for each model and find the average inter-rater agreement is around 95%. We report the average accuracy for rating and explanation across the two sets of labels, and we compute F1 score between machine-generated citation and the intersection of the two sets of manual citation. From Table 3, we see that most LLM autoevaluators obtain high rating and explanation accuracies, indicating that rating in “Yes/No” and free-form explanation generation are relatively easy tasks for the LLMs. For citation, our REC models perform the best in both completeness and faithfulness metrics, while in instruction-following, they are ranked second to llama-3.1-70B and Claude3-Opus. Specifically, for the instruction-following metric, we find REC models often cite slightly more than necessary sentences that appear to cover a conversation between a customer and an agent. In contrast, Llama-70B and Claude3 seem to be able to selectively cite the customer’s complaint and the agent’s resolution to reflect the instruction to focus on these two elements in the summary.
@@
LLM-FactVerify Wiki-FactVerify Summarization Long-form QA
Mistral-7B 68.04 56.82 55.80 62.80
Mistral-Nemo 70.18 65.27 62.14 60.88
Llama-3.1-70B 75.25 68.84 69.56 76.98
Claude-3-Opus 76.82 76.52 70.65 76.82
GPT-3.5 72.81 69.53 68.60 70.40
GPT-4 77.44 77.22 72.45 76.92
GPT-4o 76.31 79.48 72.18 75.22
REC-8B 78.04 72.46 66.27 72.73
REC-12B 78.78 85.34 72.51 77.23
REC-70B 79.67 87.48 74.04 79.12
Table 5: LLM-AggreFact performance across four common use-cases: LLM-FactVerify (ClaimVerify + FactCheck + Reveal), Wiki-FactVerify (WiCE), Summarization (AggreFact + TofuEval), and Long-form QA (ExpertQA + LFQA). The best and the second best performance for each metric are in bold and underlined separately.
@@ 6.3 General Capabilities
The RewardBench dataset assesses LLMs’ performance across several critical abilities using curated chosen-rejected response pairs Lambert et al. (2024). The evaluation involves determining the model’s win percentage based on its ability to score the “chosen” response higher than the “rejected” one. Specifically, it’s focused on testing chat performance, safety measures, and reasoning (both in terms of code and math skills), while controlling for potential biases such as overfitting to prior datasets.
We first create prompts following the RewardBench evaluation format, where each prompt consisting of a query and two potential responses. The task for LLMs is to evaluate the two responses and determine which is better. The performance of the reward models is quantified by calculating the win percentage for chosen-rejected pairs associated with each prompt. A “win” is defined by a scenario where an LLM presents a preference for the selected answer. This quantitative measure provides a clear benchmark for evaluating model performance across tasks. By averaging the results across the 4 core sections: Chat, Chat Hard, Safety, and Reasoning, a comprehensive measure of each model’s win percentage is derived. As shown in Table 4, our REC models are able to achieve SOTA performance among all generative LLMs on the RewardBench Leaderboard, demonstrating their outstanding ability to serve as an autoevaluator for pairwise comparisons.
@@
This is a benchmark for measuring the grounding capabilities of autoevaluators. Given a reference document and a claim, the autoevaluator determines if the claim is fully supported by the document. This holistic benchmark combines 10 attribution datasets used in recent studies on LLM factuality. Given some of the models such as GPT-3.5 have smaller context windows, we removed examples longer than 16K tokens from our evaluation. We use the same prompt instructions across categories, as shown in Appendix G. To reduce model API costs, we randomly sampled 256 examples per evaluation task similar to the approach used in the FLAMe Vu et al. (2024b).
Table 5 presents our attribution results on LLM-AggreFact Tang et al. (2024b), categorized into four common use-cases: (1) LLM-FactVerify: fact verification of LLM-generated responses, (2) Wiki-FactVerify: evaluating correctness of Wikipedia claims, (3) Summarization: assessing faithfulness of summaries, and (4) Long-form QA: evaluating long-form answers to questions. REC-12B and REC-70B outperform all other models in all four categories, and REC-8B outperform all the other similar-sized models (Mistral-7B and Mistral-Nemo), demonstrating their strong performance in attribution on diverse datasets.
@@ 6.4 Bias Testing
Table 6 presents the performance of the compared models on the CoBBLEr benchmark with metrics of order bias, bandwagon effect, compassion, selective bias, salience, distraction, and frequency. Note that lower scores indicates better performance  (i.e., fewer biases) on CoBBLEr. Overall, our REC-70B model has the best average performance with a score of 0.2141, followed closely by GPT-4 (0.2279) and GPT-4o (0.2349). REC-12B and REC-8B, on the other hand, outperforms all the other same-sized and smaller models (Mistral-Nemo and Mistral-7B), and even some larger models (Llama-3-70B and GPT-3.5). The results suggest that training with REC produces an LLM judge with fairer evaluation capabilities, independent of the cognitive biases tested, such as verbosity, order, or an overly authoritative tone. In contrast, off-the-shelf LLMs can be more influenced by factors such as the order of responses or the length of the text.
@@ 7 Conclusions
LLMs excel in generating coherent and high-quality text for various tasks, but evaluating content quality is challenging due to issues like factual inaccuracies and hallucination. This paper introduces three fine-tuned models—REC-8B, REC-12B, and REC-70B—designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, completeness. These models not only provide ratings for these metrics but also offer detailed explanations and verifiable citation, i.e., Ratings, Explanations, and citation (REC), thereby enhancing trust in the content. To achieve this goal, a curated dataset for explanations and citation fine-tuning is proposed to facilitate future research, which is the first public dataset containing both content quality citation and RAG citation. The models support various citation modes, including Post-fix citation, Post-fix citation mode with snippet, Inline citation, and Inline citation mode with snippet to balance the trade-off between latency and granularity of citation. Extensive evaluations on diverse benchmarks—ALCE, ExpertQA, ABCD summarization, RewardBench, LLM-AggreFact, and CoBBLEr—demonstrate that the fine-tuned LLM autoevaluator, REC-70B, outperforms state-of-the-art models. It excels in content evaluation, delivering clear explanations and ensuring citation accuracy.
@@
Our work introduces the innovative Rate, Explain, and Cite (REC) model for trusted automated AI evaluations, offering significant benefits but with a few limitations that future research will address. First, our work is limited to the baselines we compare, as non-LLM citation methods, for instance, approaches that leverage sentence embedding Reimers and Gurevych (2019) similarity to identify and cite the most semantically relevant sources were not included. Including such techniques could serve as valuable baselines to further validate our model’s generalization and effectiveness. Second, the need for post-generation explanations and citation may introduce latency. Future work will compare latencies across models and explore optimizations to improve speed. We are also considering fine-tuning the snippet output using starting and ending character numbers to optimize for latency. Finally, even though the 12B base model is capable of generating multi-lingual output, our current evaluations are limited to English datasets. Expanding to multilingual evaluations is a key goal for future research.
@@
Our contributions present a novel general-purpose autoevaluator that supports various modes of citation, as well as a dataset with citation and explanations to support further research. There are a few ethical considerations for users of such technology. As with the use of fine-tuning and RAG, it is important to be aware of potential propagating biases and feedback loops. Biased data used for training and/or retrieval could lead to model outputs that reinforces such biases, impacting model quality and fairness. Although we extensively evaluated our LLM autoevaluator and found it to be safe, unbiased, and factual, users should consider whether such considerations hold true for the input source itself. Of note, even established sources such as knowledge banks may contain inaccurate or inconsistent information. Finally, though we aimed to be transparent and consistent for our data labeling tasks, using clear instructions and validations, note that data labeling tasks are inherently subject to the bias of labeler background and experiences.